# Themis: Automatic and Efficient Deep Learning System Testing with Strong Fault Detection Capability

### v2

@@ I Introduction
Deep Learning Systems (DLSs) have been widely applied in safety-critical tasks such as autopilot and smart cities [7, 42, 35, 11, 25]. However, when a DLS (e.g., an autopilot system) is deployed in a real-world environment (e.g., a crowded city), the DLS’s input (e.g., a road condition image of the crowded city) is often perturbed by environmental noise such as raindrops and fog effects, causing incorrect values and disaster  [42, 35, 11, 25]. The incorrect output value of a DLS caused by perturbation on the DLS’s input is defined as the DLS’s fault [17].
Since a DLS’s faults greatly undermine the DLS’s reliability, a DLS must be systematically tested in order to detect as many faults as possible  [42, 35, 11, 25]. These faults are essential for further developing a DLS (i.e., improving the DLS’s accuracy by retraining the DLS with perturbed inputs which lead to faults). The rationale of existing DLS testing is inspired by conventional software testing. In conventional software testing [50, 29, 38], a testing technique generates a set of inputs (known as the test set) to explore diverse data flows of a software (e.g., statements or branches), especially the data flows which likely lead to the software’s faults such as exceptions or crashes (these data flows are called “fault-inducing data flows”). When all data flows of the software have been explored (i.e., the test coverage metric adopted by the testing technique, such as code coverage, reaches 100%), the testing process is considered completed, and the testing technique stops.
@@
We believe that the root cause behind the limitations of the existing techniques is that the triggering condition and the total number of a DLS’s fault-inducing data flows are unknown [42, 35, 17, 30, 54, 25, 11]. Hence, existing techniques inevitably require manual expert effort in inferring a set of fault-inducing data flows, in order to guide the exploration of a DLS’s data flows and the computation of test coverage (i.e., the proportion of the fault-inducing data flows being explored from the DLS during testing). Unfortunately, despite these advancements [42, 35, 17, 30, 54, 25, 11], in principle, manual effort is far away from guaranteeing that these techniques can explore a full coverage of fault-inducing data flows from a DLS, while the full coverage is an essential and sufficient condition for a testing technique to attain strong fault-detection capability [17, 30, 54]. Overall, an automated DLS testing technique that can theoretically explore a full coverage of fault-inducing data flows is highly desirable but vacant.
@@
We implemented Themis on Mindspore [41] and compared Themis against four recent and notable DLS testing techniques (Deepexplore [42], DeepGauge[35], DeepImportance [11] and Surprise Adequacy for Deep Learning System [25]), which have been deployed in the Mindspore security framework and have been evaluated by several AI developers. We evaluated Themis and these baselines on ten popular Deep Learning models (e.g., LeNet [28], VGG [48], ResNet [19]) trained on six public datasets (Cifar10 [2], ImageNet [9], Driving [52], Contagio/VirusTotal [8], Drebin [4] and MNIST [27]), which cover a complete set of datasets evaluated by the baselines [42, 35, 11, 25]. Evaluation shows that:
@@ II-C Fault-inducing data flows have to be automatically identified
Overall, these techniques allow a DLS to be tested within a reasonable time (e.g., several minutes). However, substantial studies [17, 30, 54] showed that manual effort in specifying fault-inducing data flows often causes DLS testing to have unsatisfactory fault detection capability because fault-inducing data flows are often unknown (see §I). To automatically identify fault-inducing data flows, our observation is that for any $I$ and $E(\theta)$, the $N_{i}(I+E(\theta)),1\leq i\leq n$ which induces a fault usually has a large difference with $N_{i}(I),1\leq i\leq n$ (i.e., the value of $sum_{i=1}^{n}|N_{i}(I+E(\theta))-N_{i}(I)|$ is large, we call this value as the sensitivity of a DLS’s data flow on perturbed inputs). It is because large sensitivity often results in a large difference between $M(I+E(\theta))$ and $M(I)$ (i.e., $M(I+E(\theta))$ is likely to be a fault). This observation is also confirmed in our evaluation (§VI-B). Based on this observation, we propose an input generation technique guided by sensitivity (Sensitivity Maximizing Fuzzer in §III).
@@ II-D Computing test coverage automatically
Test coverage metric measures the proportion of fault-inducing data flows explored. Ideally, when test coverage metric reaches 100%, all fault-inducing data flows of a DLS are explored [17, 30, 54]. However, since the actual number of fault-inducing data flows in a DLS is unknown, proposing an accurate test coverage metric is an open challenge [17, 30, 54]. Existing testing techniques [7, 42, 35, 11, 25] mitigate this challenge by requiring developers to specify a set of fault-inducing data flows. However, recent studies show that such manual effort is tedious and error-prone (§II-B).
@@
Figure 2: Themis’s architecture. Themis’s components are shaded in purple. The DNN’s neurons (e.g., N2) are colored in green if their MCSE value (as shown in the “Sensitivity distribution statistics”) equals 0.0 (i.e., sensitivity distribution of the neurons converged).

### v4

@@ I Introduction
Deep Learning Systems (DLSs) have been widely applied in safety-critical tasks such as autopilot and smart cities [7, 43, 36, 11, 27]. However, when a DLS (e.g., an autopilot system) is deployed in a real-world environment (e.g., a crowded city), the DLS’s input (e.g., a road condition image of the crowded city) is often perturbed by environmental noise such as raindrops and fog effects, causing incorrect values and disaster  [43, 36, 11, 27, 22, 21, 23]. The incorrect output value of a DLS caused by perturbation on the DLS’s input is defined as the DLS’s fault [16].
Since a DLS’s faults greatly undermine the DLS’s reliability, a DLS must be systematically tested in order to detect as many faults as possible  [43, 36, 11, 27]. These faults are essential for further developing a DLS (i.e., improving the DLS’s accuracy by retraining the DLS with perturbed inputs which lead to faults). The rationale of existing DLS testing is inspired by conventional software testing. In conventional software testing [51, 39], a testing technique generates a set of inputs (known as the test set) to explore diverse data flows of a software (e.g., statements or branches), especially the data flows which likely lead to the software’s faults such as exceptions or crashes (these data flows are called “fault-inducing data flows”). When all data flows of the software have been explored (i.e., the test coverage metric adopted by the testing technique, such as code coverage, reaches 100%), the testing process is considered completed, and the testing technique stops.
@@
We believe that the root cause behind the limitations of the existing techniques is that the triggering condition and the total number of a DLS’s fault-inducing data flows are unknown [43, 36, 16, 31, 27, 11]. Hence, existing techniques inevitably require manual expert effort in inferring a set of fault-inducing data flows, in order to guide the exploration of a DLS’s data flows and the computation of test coverage (i.e., the proportion of the fault-inducing data flows being explored from the DLS during testing). Unfortunately, despite these advancements [43, 36, 16, 31, 27, 11], in principle, manual effort is far away from guaranteeing that these techniques can explore a full coverage of fault-inducing data flows from a DLS, while the full coverage is an essential and sufficient condition for a testing technique to attain strong fault-detection capability [16, 31]. Overall, an automated DLS testing technique that can theoretically explore a full coverage of fault-inducing data flows is highly desirable but vacant.
@@
We implemented Themis on Mindspore and compared Themis against four recent and notable DLS testing techniques (Deepexplore [43], DeepGauge[36], DeepImportance [11] and Surprise Adequacy for Deep Learning System [27]), which have been deployed in the Mindspore security framework and have been evaluated by several AI developers. We evaluated Themis and these baselines on ten popular Deep Learning models (e.g., LeNet [30], VGG [49], ResNet [18]) trained on six public datasets (Cifar10 [2], ImageNet [9], Driving [53], Contagio/VirusTotal [8], Drebin [4] and MNIST [29]), which cover a complete set of datasets evaluated by the baselines [43, 36, 11, 27]. Evaluation shows that:
@@ II-C Fault-inducing data flows have to be automatically identified
Overall, these techniques allow a DLS to be tested within a reasonable time (e.g., several minutes). However, substantial studies [16, 31] showed that manual effort in specifying fault-inducing data flows often causes DLS testing to have unsatisfactory fault detection capability because fault-inducing data flows are often unknown (see §I). To automatically identify fault-inducing data flows, our observation is that for any $I$ and $E(\theta)$, the $N_{i}(I+E(\theta)),1\leq i\leq n$ which induces a fault usually has a large difference with $N_{i}(I),1\leq i\leq n$ (i.e., the value of $sum_{i=1}^{n}|N_{i}(I+E(\theta))-N_{i}(I)|$ is large, we call this value as the sensitivity of a DLS’s data flow on perturbed inputs). It is because large sensitivity often results in a large difference between $M(I+E(\theta))$ and $M(I)$ (i.e., $M(I+E(\theta))$ is likely to be a fault). This observation is also confirmed in our evaluation (§VI-B). Based on this observation, we propose an input generation technique guided by sensitivity (Sensitivity Maximizing Fuzzer in §III).
@@ II-D Computing test coverage automatically
Test coverage metric measures the proportion of fault-inducing data flows explored. Ideally, when test coverage metric reaches 100%, all fault-inducing data flows of a DLS are explored [16, 31]. However, since the actual number of fault-inducing data flows in a DLS is unknown, proposing an accurate test coverage metric is an open challenge [16, 31]. Existing testing techniques [7, 43, 36, 11, 27] mitigate this challenge by requiring developers to specify a set of fault-inducing data flows. However, recent studies show that such manual effort is tedious and error-prone (§II-B).
@@
Figure 2: Themis’s architecture. Themis’s components are shaded in purple. The DNN’s neurons (e.g., N2) are colored in green if their MCSE value (as shown in the “Sensitivity distribution statistics”) equals 0.0.
@@ VII Conclusion
VIII Acknowledgements
The work is supported in part by National Key R&D Program of China (2022ZD0160201), HK RIF (R7030-22), HK ITF (GHP/169/20SZ), the Huawei Flagship Research Grant in 2023, HK RGC GRF (Ref.: 17208223 and 17204424), the HKU-CAS Joint Laboratory for Intelligent System Software, and the Shanghai Artificial Intelligence Laboratory, the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Cyber Security Agency of Singapore.